If an object is projected upward with an initial velocity of 126 ft per sec, its height h after t seconds is h=-16^2+126t. Find the height of the object after 2 seconds.
Answer by stanbon(75887)

If an object is projected upward with an initial velocity of 126 ft per sec, its height h after t seconds is h=-16t^2+126t. Find the height of the object after 2 seconds.
-------------
h(2) = -16(2^2) +126(2) = -16*4 + 252 = -64+252 = 188.08 ft
-----------
Cheers,
Stan H.
---------------